OpenStudy (anonymous):

The count in a culture of bacteria was 400 after 3 hours and 25,600 after 6 hours. What is the relative rate of growth? Also, find the function that models the number of bacteria.

OpenStudy (anonymous):

Hello jah! good to see you!

OpenStudy (anonymous):

lol good to c u 2 ashes

OpenStudy (anonymous):

oops 400 after 2 hours

OpenStudy (anonymous):

hmm rate of growth is usually exponential. im tryin to remember the formula tho

OpenStudy (anonymous):

there are a couple multiple choice answers... maybe it will spark the memory

OpenStudy (anonymous):

100e^-0.0231t ; rate= 2.22s or 100e^-0.0231t; rate= 129.6 s

OpenStudy (anonymous):

general exponential equation is y=a*b^t

OpenStudy (anonymous):

hmmm or we have which looks like the correct equation n(t) = pe^(rt)

OpenStudy (anonymous):

ooh gave the wrong multiple choices earlier :( sorry

OpenStudy (anonymous):

lol i was just about to say

OpenStudy (anonymous):

would be r=1.0397; n(t)=100e^1.0397t or r=1.0397; n(t)=400e^1.0397t

OpenStudy (anonymous):

yes my apologies making you work harder and wasting your time sorry

OpenStudy (anonymous):

well the 2nd 1 doesnt make sense just by looking at it so im confident its the 1st 1

OpenStudy (anonymous):

believe it is the first one

OpenStudy (anonymous):

same

OpenStudy (anonymous):

I have a similar one if you have the time

OpenStudy (anonymous):

surei have like 10 mins

OpenStudy (anonymous):

or prefer me to repost?

OpenStudy (anonymous):

Radium-221 has a half life of 30 seconds. Find a function that models the mass remaining after t seconds given a 100g sample. Also, find how long it will take for 95% of the sample to decay.

OpenStudy (anonymous):

this goes to the multiple choice I posted first

OpenStudy (anonymous):

the 1 with 129.6 seconds

OpenStudy (anonymous):

thats how long it would take 95% to decay

OpenStudy (anonymous):

how did you come up with it if you dont mind?

OpenStudy (anonymous):

and I cant give you another medal lol

OpenStudy (anonymous):

well since the answer in the multiple choice is the same u can use that formula but for half life i like the formula: Amount remaining=(original amount)/(2^(t/decay time)) so u just plug it in and solve 95% is gone so 5% remain or 5 grams since its out of 100 5=100/(2^(t/30))

OpenStudy (anonymous):

solve for t and its ok lol i dont really do this for the medals anyway :)

OpenStudy (anonymous):

oooh that helps a ton... thank you. Somehow I missed the other question so will rerwork that

OpenStudy (anonymous):

the 1 before was wrong?

OpenStudy (anonymous):

yes... not sure why

OpenStudy (anonymous):

the only difference is the 100 and the 400 right? which did u choose?

OpenStudy (anonymous):

the 1 with 100 should b right... hmm mayb re-post and ask again